OpenStudy (anonymous):

Calculus 1 Inverse question!

OpenStudy (anonymous):

if \[f ^{-1}\] is the inverse of f, determine the value of \[f(f ^{-1}(3))\]

OpenStudy (anonymous):

Answer Choices: 1. Need to know f 2. f(f−1(3)) =1/3 3. f(f−1(3)) =1/9 4. f(f−1(3)) = 3 5. f(f−1(3)) = 9

OpenStudy (psymon):

It's always going to return the same value. You take the inverse at a point and then plug that back into the original function, you just get the same point in return. So it'd be 3.

OpenStudy (anonymous):

@Psymon I just checked and that answer isn't correct

OpenStudy (psymon):

Interesting. The only reason I even said that was because I physically made examples and tried them out.

OpenStudy (psymon):

And the way I think of it, it's just like how it works with trig functions. \[\sin(\sin^{-1} x) = x\] \[\sin^{-1} (sinx)=x \]

OpenStudy (anonymous):

@Psymon you're answer was actually correct. for some reason the website I'm using counted it wrong.

OpenStudy (anonymous):

sorry about that.

OpenStudy (psymon):

Hm O.o Interesting. Happens unfortunately. Glad you were able to see the correct answer anyway ^_^